The weights of the students in a certain year level are normally distributed with a mean of 60kg and a standard deviation of 3.5kg. Find the probability that a student randomly selected from this group weighs. i) 55kg or less. ii) 68kg or more. iii) between 48kg and 53kg




1
Expert's answer
2021-02-24T07:58:38-0500

The weight of the students in a certain year level are normally distributed with mean "60kg" and standard deviation "3.5kg" .

Let "X" be random variable which defines weight of the students.

Then , "\mu =60" and "\sigma =3.5"

Let us take "Z=\frac{X-\mu}{\sigma }" . Then "Z=\frac{X-60}{3.5}"

(i) Probability that a student randomly selected is less than or equal to "55" kg is "P(X\leq 55)" .

"\therefore P(X\leq 55)=P(Z \leq \frac{55-60}{3.5})"

"=P(Z \leq -1.43)"

"=0.5-P(0<Z<1.43)"

"=0.5-0.4236" [ from normal distribution table]

"=0.076" (approximately)

(ii) Probability that a student randomly selected is greater than or equal to "68" kg is "P(X\geq 68)" .

"\therefore P(X\geq 68)=P(Z \geq \frac{68-60}{3.5})"

"=P(Z\geq 2.29)"

"=0.5-P(0<Z<2.29)"

"=0.5-0.4889"

"=0.011" (approximately)

(iii)Probability that a student randomly selected is greater than "48" kg and less than 53 kg is "P(48<X<53)"

"\therefore P(48<X<53)=P(\frac{48-60}{3.5}<Z<\frac{53-60}{3.5})"

"=P(-3.43<Z<-2)"

"=P(0<Z<3.43)-P(0<Z<2)"

"=0.4996-0.4772"

"=0.022" (approximately)